OpenStudy (anonymous):

Multiple Response: Please select all correct answers and click "submit." Which of the following could be the equation of the graph shown below? Check all that apply. A. y = -4x + 2 B. 3x + 2y = 6 C. y = 5 D. 2x - y = 7

OpenStudy (anonymous):

@johnweldon1993

Directrix (directrix):

As you look at the line from left to right, the line is falling. That is, as x gets larger, y gets smaller. That means the slope is negative. Look at the options and eliminate those which have positive slopes.

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (owlcoffee):

Just to add Directrix's explaination, lines have this element called "slope" which signifies the angle in regards to the x-axis. When the slope is positive the line looks like this:  And when it is negative:  The slope can be put to evidence in a algebraic form by using the form: \[y=mx+c \] This is the "point-slope" form of the line and can model each line I showed you above, "m" is the angular coefficient, a more fancy way to say "slope".

OpenStudy (anonymous):

@narissa

OpenStudy (mrnood):

in addition: the value of c is the point at which the graph intersects the y axis in your graph this is a positive number so - look a the equations re-arrange them into y=mx+b format choose those that have a negative slope AND a positive intersect

OpenStudy (anonymous):

at the end what are the final answers

OpenStudy (anonymous):

because i did it but i got that question wrong

OpenStudy (driftracer305):

@selena929....it should be A

OpenStudy (anonymous):

but that was the only answer its multi choice

OpenStudy (anonymous):

@DavidUsa what do you think ?

OpenStudy (davidusa):

estoy estudiando para el amc 8.

OpenStudy (mrnood):

A. y = -4x + 2 B. 3x + 2y = 6 2y =-3x+6 y=-3/2 x+3 C. y = 5 D. 2x - y = 7 y=2x-7 look for negative slope m AND positive intersect b